OpenStudy (safyousuff97):

Anyone willing to help me on these problems? (not good in math so i need step by step help , i have know idea what to do )

OpenStudy (1andrewalarcon1):

Ill give it a go

OpenStudy (safyousuff97):

first one Task 1 Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (1andrewalarcon1):

Sorry, Ill be honest and say i am also confused by this...

OpenStudy (safyousuff97):

thats ok :) thanks for trying to help tho

OpenStudy (whpalmer4):

okay, first part: \[f(x) = x+a\]\[g(x) = cx-d\] The way you find the inverse of \[f(x) = \text{<something>}\]is to replace \(f(x)\) with \(y\), and then go through and swap \(x\) and \(y\). solve for \(y\). that's your inverse. So, say you need to find the inverse of \[f(x) = 2x+4\] replace \(f(x)\) with \(y)\:\[y=2x+4\]now swap \(x\) and \(y\): \[x = 2y + 4\]solve for \(y\): \[x-4 = 2y+4-4\]\[x-4=2y\]\[\frac{x-4}{2} = \frac{2y}2\] \[y=\frac{x}{2}-2\]

OpenStudy (whpalmer4):

so, how do we find the inverse of \[f(x) = x+a\]given what I just told you?

OpenStudy (safyousuff97):

oh nvm, how do we find an inverse?

OpenStudy (21younga):

okay so first do f(x)=3x-5 okay? well to get g(x) you find the inverse of f(x) like this: f(x)= 3x -5 Now change f(x) for y=3x-5 switch the x and y x= 3y-5 Now isolate the y by adding 5 to both sides x+5=3y then divide both sides by 5 x=5=3y/5 y=x+5/3 For part one you write the functions, and then for part 2, you show the work that i did to find the inverse. Does this make sense? I found this on another Open-study site that's how i got all of this info :)

OpenStudy (safyousuff97):

yes thank u so much

OpenStudy (21younga):

Welcome :)

OpenStudy (safyousuff97):

can u help with the rest?

OpenStudy (safyousuff97):

@mathwizzard3

OpenStudy (mathwizzard3):

I might be able to help....with which part?

OpenStudy (mathwizzard3):

I'll ask my buddy to come over to maybe help @chalkskull

OpenStudy (safyousuff97):

the last part thanks

OpenStudy (mathwizzard3):

So what have you come up with so far?

OpenStudy (mathwizzard3):

aka your two functions

OpenStudy (safyousuff97):

i don't understand what the others said for part one and part 2 so i don't know what i should do for the last part

OpenStudy (mathwizzard3):

Yeah.....it seems pretty confusing I'll ask some people to help @chalkskull @alibaby @milk333 @matlee @mathstudent55

OpenStudy (safyousuff97):

ok

OpenStudy (mathwizzard3):

@agent0smith @AnitaS. @TheCuteOne25 @HarleyQuinn98 @Thebossofme191

OpenStudy (qwertty123):

Do you still need help? @safyousuff97

OpenStudy (jdoe0001):

@safyousuff97 so. hmmm is it \(\bf f(x)=\cfrac{x+a}{b}\qquad g(x)=cx-d\)

OpenStudy (safyousuff97):

yes please thank you

OpenStudy (qwertty123):

With what?